# Moment bounds for condition numbers and singular values of high-dimensional Gaussian random matrices: Applications and limitations

### v1

@@ 3.1 Prediction Risk for ridge regression with random predictors
Consider the random-design multi-response linear model in (23) and define $R_{\mathrm{pred}}(\tilde{\lambda}_{n}\mid X_{n})$ as in (24). Assume that the rows of $X_{n}\in\mathbb{R}^{n\times p}$ are independent and identically distributed as $N(0,\tilde{\Sigma}_{n})$, where $\tilde{\Sigma}_{n}$ satisfies (1). Assume further that the rows of $E$ are independent and identically distributed as $N(0,\Sigma_{\varepsilon})$, where $\Sigma_{\varepsilon}$ satisfies (1). Suppose that $pq/n\to 0$ as $n\to\infty$, and fix $\lambda>0$.
Then there exists a constant $K>0$ (depending only on $\lambda,c_{m},c_{M},c_{\varepsilon,m},c_{\varepsilon,M}$) such that, for all sufficiently large $n$,
@@
In particular, if $\|B\|_{F}/n\to 0$, then $\mathbb{E}\big[R_{\mathrm{pred}}(\tilde{\lambda}_{n}\mid X_{n})\big]\to 0$ as $n\to\infty$.
@@
In [14, Proposition 2], the author establish an almost sure limit for $R_{\mathrm{pred}}(\tilde{\lambda}_{n}\mid X_{n})$ in the case $q=1$ under the regime $p/n\to\gamma\in[0,1)$, while imposing substantially weaker assumptions than Gaussianity. Specifically, in our notation they consider the design $X_{n}=Z_{n}\tilde{\Sigma}_{n}^{1/2},$ where $Z_{n}$ is an $n\times p$ random matrix with i.i.d. entries having zero mean, unit variance, and a finite fourth moment, and where $\tilde{\Sigma}_{n}$ satisfies (1). Although this assumption is weaker than the Gaussian design we adopt to prove Theorem 6, it is important to emphasize that our objective is different. Rather than almost sure convergence, we require convergence of the mean prediction risk, i.e., after taking expectation over the randomness of $X_{n}$. This constitutes a stronger requirement than the almost sure statement in [14, Proposition 2].
Indeed, removing the Gaussianity assumption from Theorem 6 is not always possible, as we demonstrate in Section 4. In particular, we construct a sub-Gaussian counterexample (see Proposition 11) satisfying all the assumptions of [14, Proposition 2], yet for which the expectation of the inverse of the (nonzero) minimum singular value diverges to infinity, a quantity that plays a key role in bounding the expected prediction risk. At the same time, the corresponding almost sure convergence may still hold by the Bai–Yin law [4], yielding a result consistent with [14, Proposition 2].
@@ 3.2 Expected estimation risk for inverse covariance matrices
Another widely used Gram matrix in statistical analysis is the sample variance–covariance matrix. Let $X_{1},\ldots,X_{n}\in\mathbb{R}^{p}$ be i.i.d. centered Gaussian random vectors with covariance matrix $\Sigma$, and define the sample covariance matrix
$S_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}.$
@@
$\big(\mathbb{E}\|S_{n}-\Sigma\|_{2}^{r}\big)^{1/r}=O\Big(\sqrt{\frac{p}{n}}\vee\frac{p}{n}\Big),$
@@
In the next theorem, we show that under Gaussianity and the same bounded eigenvalue assumptions as in (1) on $\Sigma$, for any fixed $r\geq 1$ the inverse covariance estimation error i.e. $\mathbb{E}[\|S_{n}^{-1}-\Sigma^{-1}\|_{2}^{r}]$ achieves the same rate as the corresponding moment bound for $\mathbb{E}[\|S_{n}-\Sigma\|_{2}^{r}]$. This yields a strong stability achieves for inverse covariance estimation in the Gaussian case. In the subsequent section (see Proposition 12), we show that this behavior is specific to Gaussian designs and, unlike the moment bounds for $\|S_{n}-\Sigma\|_{2}$, need not hold for general sub-Gaussian distributions.
@@
Let $X_{1},\ldots,X_{n}\in\mathbb{R}^{p}$ be i.i.d. centered Gaussian random vectors with covariance matrix $\Sigma$, and define
$S_{n}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}.$
@@
Write $X_{i}=\Sigma^{1/2}Z_{i}$ with $Z_{i}\sim N(0,I_{p})$ i.i.d., and let $Z\in\mathbb{R}^{n\times p}$ be the matrix with rows $Z_{i}^{\top}$. Then
@@ 3.3 Solving Gram linear systems: Expected iteration complexity and the role of condition number moments
Systems of the form (31) arise frequently in statistics and machine learning, for example through normal equations, iterative least-squares subproblems, and kernel methods. In such settings, the condition number plays a fundamental role because the righthand side vector $b$ typically depends on the observed data. Consequently, it is important to understand how perturbations or noise in the data, which directly induce perturbations in $b$, affect the optimal solution. In particular, as discussed by [23], the condition number $\kappa(X_{n})$ governs the sensitivity of the solution to changes in $b$, and hence directly influences stability and generalization. A representative example comes from supervised learning, where the goal is to estimate an unknown function from measurements at random points in a high-dimensional space, and where low sensitivity to errors in the data is essential for good predictive performance. In this section, we show that studying condition numbers is therefore unavoidable for understanding the computational behavior and convergence properties of algorithms used to solve such Gram-type linear systems. In the growing-dimensional regime where $p/n\to\gamma\in[0,1]$ and $p\leq n$, the Gram matrix satisfies $S_{n}\succ 0$. Consequently, the linear system in (31) admits a unique solution given by
$\theta_{n}^{\star}=S_{n}^{-1}b.$
In contrast, when $p/n\to\gamma\in[1,\infty)$ and $p>n$, the system may fail to admit a solution unless $b\in\operatorname{range}(X_{n}^{\top})$ (or equivalently $\operatorname{range}(S_{n})$). This condition holds in many statistical settings—for instance, in least-squares normal equations arising from linear regression. Even when a solution exists, it is generally not unique. Among all solutions, to obtain a unique minimum-$\ell_{2}$-norm solution, we define $\theta_{n}^{\star}=S_{n}^{+}b$, where $S_{n}^{+}$ denotes the Moore–Penrose pseudo-inverse of $S_{n}$ [21].
@@
An iterative gradient descent approach Although $\theta_{n}^{\star}$ admits a closed-form representation, computing it directly becomes impractical when $p$ is large, since it requires forming and factorizing the $p\times p$ matrix $S_{n}$. This motivates the use of iterative methods-under the assumption $b\in\operatorname{range}(X_{n}^{\top})$—which avoid explicit inversion and instead exploit matrix-vector products; see, for example, [27, 32].
@@
We will now show that, under the Gaussianity assumption and for sufficiently large sample size $n$, and under an appropriate choice of step size $\eta$, the expected iteration complexity of gradient descent for solving the Gram system (31) remains bounded by a constant independent of $n$ whenever $p/n\to\gamma\in[0,\infty)\setminus\{1\}$ (see Theorem 9). The key observation is that, even though $g_{n}$ may fail to be strongly convex on $\mathbb{R}^{p}$ when $S_{n}$ is rank-deficient, Lemma 8 guarantees linear convergence of the iterates on $\operatorname{range}(S_{n})$, provided that $\theta^{(0)}\in\operatorname{range}(S_{n})$ and $\eta$ is chosen appropriately. As a result, the iteration complexity is governed by the square of the condition number of $X_{n}$. In particular, controlling the second moment of $\kappa(X_{n})$ is central to obtaining uniform bounds on the expected runtime.
@@
$\mathbb{E}\big[T_{\varepsilon,n}\big]\leq C\log\!\left(\frac{1}{\varepsilon}\right).$
@@
First, note that (1) implies $\tilde{\Sigma}_{n}\succ 0$ for all $n$, and therefore $\mathrm{rank}(\tilde{X}_{n})=\min(n,p)$ almost surely. In particular, in the wide-matrix regime $p>n$, the smallest nonzero eigenvalue of $S_{n}$ equals the square of the smallest nonzero singular value of $X_{n}$ (equivalently, the $n$-th singular value).
@@
since each iterate can be written as a polynomial in $S_{n}$ applied to $b$; see, for example, [27, 32]. More generally, Krylov subspace methods (e.g., the conjugate gradient algorithm) construct iterates by optimizing over these subspaces, often yielding faster convergence for symmetric positive definite systems such as (31); see, for example, the discussion after [32, Theorem 38.5]. In particular, such methods ensure $T_{\varepsilon,n}=O(\kappa(X_{n}))$. However, even in this case Theorem 1 and Lemma 2 (with $r=1$) will ensure that $\mathbb{E}\big[T_{\varepsilon,n}\big]$ is finite when $p/n\to\gamma\in[0,\infty)\setminus\{1\}$.
@@
Assume that the rows of $X_{n}\in\mathbb{R}^{n\times p}$ are independent and identically distributed as $N(0,\tilde{\Sigma}_{n})$, where $\tilde{\Sigma}_{n}$ satisfies (1). Fix $\varepsilon\in(0,1)$, define the worst case version of $T_{\varepsilon,n}$
@@
Finally, if $p/n\to\gamma=1$ for $n\to\infty$ Then, for every fixed $\varepsilon\in(0,1)$,
@@ 4 Sub-Gaussian Extensions: Limitations and Caveats
Let $X\in\mathbb{R}^{n\times p}$ have i.i.d. entries $\{X_{ij}\}$ supported on $(0,1)$ with distribution function
$F(u)=\frac{1}{\log\!\big(\frac{e}{u}\big)},\qquad u\in(0,1).$
Then $X_{ij}$ is sub-Gaussian (hence $X$ has i.i.d. sub-Gaussian entries). Suppose $\lim_{n\rightarrow\infty}\frac{p}{n}=:\gamma$. Then
$\mathbb{E}\!\left[\frac{\sqrt{n}}{s_{\min}(X)}\right]\to\infty,\qquad\mathbb{E}\big[\kappa(X)\big]\to\infty,$
for any $\gamma\in[0,\infty)$.
@@
Since $0<X_{ij}<1$ almost surely, the random variable $X_{ij}$ is bounded and therefore sub-Gaussian.
@@
$E_{j}(\epsilon)=\{X_{ij}\leq\epsilon\text{ for all }i=1,\dots,n\}.$
By independence,
@@
$\mathbb{P}\!\left(\frac{\sqrt{n}}{s_{\min}(X)}\geq\frac{1}{\epsilon}\right)\geq\left(\frac{1}{\log\!\big(\frac{e}{\epsilon}\big)}\right)^{n}.$ (35)
@@
$\mathbb{E}\!\left[\frac{\sqrt{n}}{s_{\min}(X)}\right]$ $\geq\int_{1}^{\infty}\mathbb{P}\!\left(\frac{\sqrt{n}}{s_{\min}(X)}\geq t\right)\,dt\geq\int_{1}^{\infty}\left(\frac{1}{\log(et)}\right)^{n}dt.$
The last integral diverges for every fixed $n\geq 1$ since $\int_{1}^{\infty}(\log t)^{-n}\,dt\to\infty$ as $n\to\infty$. This proves $\mathbb{E}[\sqrt{n}/s_{\min}(X)]\to\infty$ as $n\to\infty$.
To show $\mathbb{E}[\kappa(X)]\to\infty$, drop column $j$ and denote the remaining matrix by $X_{-j}$. Since the entries are i.i.d. and independent across columns, the event $E_{j}(\epsilon)$ is independent of any event determined by $X_{-j}$. Let $m_{2}=\mathbb{E}[X_{ij}^{2}]\in(0,1)$ and define
@@
Because $0\leq X_{ij}^{2}\leq 1$, a Bernstein-type inequality ([33]) yields $\mathbb{P}(G_{j})\geq 1-\exp\{-cn(p-1)\}$ for some constant $c>0$, and in particular $\mathbb{P}(G_{j})\geq 1/2$ for all sufficiently large $n$. Set $Q_{0}(n):=\mathbb{P}(G_{j})$.
@@
$s_{\max}(X)\geq\frac{\|X\|_{F}}{\sqrt{p}}\geq\frac{\|X_{-j}\|_{F}}{\sqrt{p}}\geq\left(\sqrt{\frac{m_{2}}{4}\frac{p-1}{p}}\right)\sqrt{n}=:C^{*}(p)\sqrt{n}.$
@@
$\kappa(X)=\frac{s_{\max}(X)}{s_{\min}(X)}\geq\frac{C^{*}(p)\sqrt{n}}{\epsilon\sqrt{n}}=\frac{C^{*}(p)}{\epsilon}.$
@@
$\mathbb{P}\!\left(\kappa(X)\geq\frac{C^{*}(p)}{\epsilon}\right)\geq\mathbb{P}(E_{j}(\epsilon))\,\mathbb{P}(G_{j})=Q_{0}(n)\left(\frac{1}{\log\!\big(\frac{e}{\epsilon}\big)}\right)^{n}.$
@@
$\mathbb{P}\big(\kappa(X)\geq t\big)\geq Q_{0}(n)\left(\frac{1}{\log\!\big(\frac{et}{C^{*}(p)}\big)}\right)^{n}.$
@@
$\mathbb{E}[\kappa(X)]\geq\int_{C^{*}(p)}^{\infty}\mathbb{P}(\kappa(X)\geq t)\,dt\geq\frac{1}{2}\int_{\sqrt{\frac{m_{2}}{4}}}^{\infty}\left(\frac{1}{\log\!\big(\frac{8et}{\sqrt{m_{2}}}\big)}\right)^{n}dt,$
for sufficiently large $n$. The last step follows since for all sufficiently large $n$, $Q_{0}(n)\geq 1/2$, while $C^{*}(p)\in[1/2,1)$ for every $p\geq 2$. Hence, the last integral diverges, implying $\mathbb{E}[\kappa(X)]\to\infty$ as $n\to\infty$. This completes the proof.
@@
Proposition 11 immediately implies that for every $r>1$ and $n\to\infty$,
$\mathbb{E}\!\left[\left(\frac{\sqrt{n}}{s_{\min}(X)}\right)^{r}\right]\to\infty,\qquad\mathbb{E}\!\left[\kappa(X)^{r}\right]\to\infty,$

### v2

@@ 3.1 Prediction Risk for ridge regression with random predictors
Consider the random-design multi-response linear model in (23) and define $R_{\mathrm{pred}}(\tilde{\lambda}_{n}\mid X_{n})$ as in (24). Assume that the rows of $X_{n}\in\mathbb{R}^{n\times p}$ are independent and identically distributed as $N(0,\tilde{\Sigma}_{n})$, where $\tilde{\Sigma}_{n}$ satisfies (1). Assume further that the rows of $E_{n}$ are independent and identically distributed as $N(0,\Sigma_{\varepsilon})$, where $\Sigma_{\varepsilon}$ satisfies (1) with corresponding lower bound $c_{\varepsilon,m}$ and upper bound $c_{\varepsilon,M}$.
Then for a fixed $\lambda>0$ there exists a constant $K>0$ (depending only on $\lambda,c_{m},c_{M},c_{\varepsilon,m},c_{\varepsilon,M}$) such that, for all sufficiently large $n$,
@@
In particular, if $\|B\|_{F}/n\to 0$ and $pq/n\to 0$ as $n\to\infty$, then $\mathbb{E}\big[R_{\mathrm{pred}}(\tilde{\lambda}_{n}\mid X_{n})\big]\to 0$ as $n\to\infty$.
@@
In [14, Proposition 2], the authors establish an almost sure limit for $R_{\mathrm{pred}}(\tilde{\lambda}_{n}\mid X_{n})$ in the case $q=1$ under the regime $p/n\to\gamma\in[0,1)$, while imposing substantially weaker assumptions than Gaussianity. Specifically, in our notation they consider the design $X_{n}=Z_{n}\tilde{\Sigma}_{n}^{1/2},$ where $Z_{n}$ is an $n\times p$ random matrix with i.i.d. entries having zero mean, unit variance, and a finite fourth moment, and where $\tilde{\Sigma}_{n}$ satisfies (1). Although this assumption is weaker than the Gaussian design we adopt to prove Theorem 6, it is important to emphasize that our objective is different. Rather than almost sure convergence, we require convergence of the mean prediction risk, i.e., after taking expectation over the randomness of $X_{n}$. This constitutes a stronger requirement than the almost sure statement in [14, Proposition 2].
Indeed, removing the Gaussianity assumption from Theorem 6 is not always possible, as we demonstrate in Section 4. In particular, we construct a sub-Gaussian counterexample (see Proposition 11) satisfying all the assumptions of [14, Proposition 2], yet for which the expectation of the inverse of the minimum singular value diverges to infinity, a quantity that plays a key role in bounding the expected prediction risk. At the same time, the corresponding almost sure convergence may still hold by the Bai–Yin law [4], yielding a result consistent with [14, Proposition 2].
@@ 3.2 Expected estimation risk for inverse covariance matrices
Another widely used Gram matrix in statistical analysis is the sample variance–covariance matrix. Let $Y_{1},\ldots,Y_{n}\in\mathbb{R}^{p}$ be i.i.d. centered Gaussian random vectors with covariance matrix $\Sigma$, and define the sample covariance matrix
$S_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}Y_{i}^{\top}.$
@@
$\big(\mathbb{E}\|S_{n}-\Sigma\|_{2}^{r}\big)^{1/r}=O\Big(\sqrt{\frac{p}{n}}\vee\frac{p}{n}\Big),$ (30)
@@
In the next theorem, we show that under Gaussianity and the same bounded eigenvalue assumptions as in (1) on $\Sigma$, for any fixed $r\geq 1$ the inverse covariance estimation error i.e. $\mathbb{E}[\|S_{n}^{-1}-\Sigma^{-1}\|_{2}^{r}]$ achieves the same rate as the corresponding moment bound for $\mathbb{E}[\|S_{n}-\Sigma\|_{2}^{r}]$. This yields a strong stability for inverse covariance estimation in the Gaussian case. In the subsequent section (see Proposition 12), we show that this behavior is specific to Gaussian designs([15]) and, unlike the moment bounds for $\|S_{n}-\Sigma\|_{2}$, need not hold for general sub-Gaussian distributions.
@@
Let $Y_{1},\ldots,Y_{n}\in\mathbb{R}^{p}$ be i.i.d. centered Gaussian random vectors with covariance matrix $\Sigma$, and define
$S_{n}:=\frac{1}{n}\sum_{i=1}^{n}Y_{i}Y_{i}^{\top}.$
@@
Write $Y_{i}=\Sigma^{1/2}Z_{i}$ with $Z_{i}\sim N(0,I_{p})$ i.i.d., and let $Z\in\mathbb{R}^{n\times p}$ be the matrix with rows $Z_{i}^{\top}$. Then
@@ 3.3 Solving Gram linear systems: Expected iteration complexity and the role of condition number moments
Systems of the form (32) arise frequently in statistics and machine learning, for example through normal equations, iterative least-squares subproblems, and kernel methods. In such settings, the condition number plays a fundamental role because the righthand side vector $b$ typically depends on the observed data. Consequently, it is important to understand how perturbations or noise in the data, which directly induce perturbations in $b$, affect the optimal solution. In particular, as discussed by [23], the condition number $\kappa(X_{n})$ governs the sensitivity of the solution to changes in $b$, and hence directly influences stability and generalization. A representative example comes from supervised learning, where the goal is to estimate an unknown function from measurements at random points in a high-dimensional space, and where low sensitivity to errors in the data is essential for good predictive performance. In this section, we show that studying condition numbers is therefore unavoidable for understanding the computational behavior and convergence properties of algorithms used to solve such Gram-type linear systems. In the growing-dimensional regime where $p/n\to\gamma\in[0,1]$ and $p\leq n$, the Gram matrix satisfies $S_{n}\succ 0$. Consequently, the linear system in (32) admits a unique solution given by $\theta_{n}^{\star}=S_{n}^{-1}b.$
In contrast, when $p/n\to\gamma\in[1,\infty)$ and $p>n$, the system may fail to admit a solution unless $b\in\operatorname{range}(X_{n}^{\top})$ (or equivalently $\operatorname{range}(S_{n})$). This condition holds in many statistical settings—for instance, in least-squares normal equations arising from linear regression. Even when a solution exists, it is generally not unique. Among all solutions, to obtain a unique minimum-$\ell_{2}$-norm solution, we define $\theta_{n}^{\star}=S_{n}^{+}b$, where $S_{n}^{+}$ denotes the (unique) Moore–Penrose pseudo-inverse of $S_{n}$ [21]. To summarize, the definition of $\theta_{n}^{\star}$ across dimensional regimes can be written as
$\theta_{n}^{\star}=\begin{cases}S_{n}^{-1}b,&\text{if }p\leq n,\\[8.0pt] S_{n}^{+}b,&\text{if }p>n.\end{cases}$
@@
An iterative gradient descent approach Although $\theta_{n}^{\star}$ admits a closed-form representation, computing it directly becomes impractical when $p$ is large, since it requires forming and factorizing the $p\times p$ matrix $S_{n}$. This motivates the use of iterative methods–under the assumption $b\in\operatorname{range}(X_{n}^{\top})$–which avoid explicit inversion and instead exploit matrix-vector products; see, for example, [26, 30].
@@
We will now show that, under the Gaussianity assumption and for sufficiently large sample size $n$, and under an appropriate choice of step size $\eta$, the expected iteration complexity of gradient descent for solving the Gram system (32) remains bounded by a constant independent of $n$ whenever $p/n\to\gamma\in[0,\infty)\setminus\{1\}$. The key observation is that, even though $g_{n}$ may fail to be strongly convex on $\mathbb{R}^{p}$ when $S_{n}$ is rank-deficient, Lemma 8 guarantees linear convergence of the iterates on $\operatorname{range}(S_{n})$, provided that $\theta^{(0)}\in\operatorname{range}(S_{n})$ and $\eta$ is chosen appropriately. As a result, the iteration complexity is governed by the square of the condition number of $X_{n}$. In particular, controlling the second moment of $\kappa(X_{n})$ is central to obtaining uniform bounds on the expected runtime.
@@
$\mathbb{E}\big[T_{\varepsilon,n}\big]\leq C\log\!\left(\frac{1}{\varepsilon}\right)+1.$
@@
First, note that (1) implies $\tilde{\Sigma}_{n}\succ 0$ for all $n$, and therefore $\mathrm{rank}(\tilde{X}_{n})=\min(n,p)$ almost surely. In particular, in the wide-matrix regime $p>n$, the smallest nonzero eigenvalue of $S_{n}$ equals the square of the smallest singular value of $X_{n}$ (equivalently, the $n$-th singular value).
@@
since each iterate can be written as a polynomial in $S_{n}$ applied to $b$; see, for example, [26, 30]. More generally, Krylov subspace methods (e.g., the conjugate gradient algorithm) construct iterates by optimizing over these subspaces, often yielding faster convergence for symmetric positive definite systems such as (32); see, for example, the discussion after [30, Theorem 38.5]. In particular, such methods ensure $T_{\varepsilon,n}=O(\kappa(X_{n}))$. However, even in this case Theorem 1 and Lemma 2 (with $r=1$) will ensure that $\mathbb{E}\big[T_{\varepsilon,n}\big]$ is finite when $p/n\to\gamma\in[0,\infty)\setminus\{1\}$.
@@
Assume that the rows of $X_{n}\in\mathbb{R}^{n\times p}$ are independent and identically distributed as $N(0,\tilde{\Sigma}_{n})$, where $\tilde{\Sigma}_{n}$ satisfies (1). Fix $\varepsilon\in(0,1)$, and define the worst case version of $T_{\varepsilon,n}$
@@
Finally, if $p/n\to\gamma=1$ for $n\to\infty$, then, for every fixed $\varepsilon\in(0,1)$,
@@ 4 Sub-Gaussian Extensions: Limitations and Caveats
Let $X_{n}\in\mathbb{R}^{n\times p}$ have i.i.d. entries $\{X_{ij}\}$, where $X_{ij}=Z_{ij}U_{ij}$ for all $(i,j)$. The variables $U_{ij}$ (supported on (0,1)) are i.i.d. with distribution function
$F(u)=\frac{1}{\log\!\big(\frac{e}{u}\big)},\qquad u\in(0,1).$ (36)
The variables $Z_{ij}$ are i.i.d. Rademacher random variables independent of $U_{ij}$; that is, $P(Z_{ij}=1)=P(Z_{ij}=-1)=\frac{1}{2}$. Then $X_{ij}$ is mean-zero sub-Gaussian (hence $X_{n}$ has i.i.d. mean-zero sub-Gaussian entries). Then, for every $n\geq 1$ and $p\geq 2$,
$\mathbb{E}\!\left[\frac{\sqrt{n}}{s_{\min}(X_{n})}\right]=\infty,\qquad\mathbb{E}\big[\kappa(X_{n})\big]=\infty.$
@@
Since $1<X_{ij}<1$ almost surely, the random variable $X_{ij}$ is bounded and therefore sub-Gaussian. Also, $\mathbb{E}[X_{ij}]=\mathbb{E}[Z_{ij}U_{ij}]=\mathbb{E}[Z_{ij}]\mathbb{E}[U_{ij}]=0$.
@@
$E_{j}(\epsilon)=\{|X_{ij}|\leq\epsilon\text{ for all }i=1,\dots,n\}.$
By independence and since $|X_{ij}|\overset{d}{=}U_{ij}$,
@@
$\mathbb{P}\!\left(\frac{\sqrt{n}}{s_{\min}(X_{n})}\geq\frac{1}{\epsilon}\right)\geq\left(\frac{1}{\log\!\big(\frac{e}{\epsilon}\big)}\right)^{n}.$ (37)
@@
$\mathbb{E}\!\left[\frac{\sqrt{n}}{s_{\min}(X_{n})}\right]$ $\geq\int_{1}^{\infty}\mathbb{P}\!\left(\frac{\sqrt{n}}{s_{\min}(X_{n})}\geq t\right)\,dt\geq\int_{1}^{\infty}\left(\frac{1}{\log(et)}\right)^{n}dt.$
The last integral is infinite for every fixed $n\geq 1$ since $\int_{1}^{\infty}(\log t)^{-n}\,dt=\infty$ for every fixed $n\geq 1$. This proves $\mathbb{E}[\sqrt{n}/s_{\min}(X_{n})]=\infty$ for every fixed $n\geq 1$.
To show $\mathbb{E}[\kappa(X_{n})]=\infty$, drop column $j$ from matrix $X_{n}$ and denote the remaining matrix by $X_{-j}$. Since the entries are i.i.d. and hence independent across columns, the event $E_{j}(\epsilon)$ is independent of any event determined by $X_{-j}$. Let $m_{2}=\mathbb{E}[X_{ij}^{2}]\in(0,1)$ and define
@@
Because $0\leq X_{ij}^{2}\leq 1$, a Bernstein-type inequality ([31]) yields $\mathbb{P}(G_{j})\geq 1-\exp\{-cn(p-1)\}$ for some universal constant $c>0$. Set $Q_{0}(n):=\mathbb{P}(G_{j})$ and note that $Q_{0}(n)>0$ for every $n\geq 1$ and $p\geq 2$.
@@
$s_{\max}(X_{n})\geq\frac{\|X_{n}\|_{F}}{\sqrt{p}}\geq\frac{\|X_{-j}\|_{F}}{\sqrt{p}}\geq\left(\sqrt{\frac{m_{2}}{4}\frac{p-1}{p}}\right)\sqrt{n}=:C^{*}(p)\sqrt{n}.$
@@
$\kappa(X_{n})=\frac{s_{\max}(X_{n})}{s_{\min}(X_{n})}\geq\frac{C^{*}(p)\sqrt{n}}{\epsilon\sqrt{n}}=\frac{C^{*}(p)}{\epsilon}.$
@@
$\mathbb{P}\!\left(\kappa(X_{n})\geq\frac{C^{*}(p)}{\epsilon}\right)\geq\mathbb{P}(E_{j}(\epsilon))\,\mathbb{P}(G_{j})=Q_{0}(n)\left(\frac{1}{\log\!\big(\frac{e}{\epsilon}\big)}\right)^{n}.$
@@
$\mathbb{P}\big(\kappa(X_{n})\geq t\big)\geq Q_{0}(n)\left(\frac{1}{\log\!\big(\frac{et}{C^{*}(p)}\big)}\right)^{n}.$
@@
$\mathbb{E}[\kappa(X_{n})]\geq\int_{C^{*}(p)}^{\infty}\mathbb{P}(\kappa(X_{n})\geq t)\,dt\geq Q_{0}(n)\int_{\sqrt{\frac{m_{2}}{4}}}^{\infty}\left(\frac{1}{\log\!\big(\frac{8et}{\sqrt{m_{2}}}\big)}\right)^{n}dt,$
for $n\geq 1$. The last step follows since $(p-1)/p\in[1/2,1)$ for every $p\geq 2$. Hence, the last integral diverges since for all $n\geq 1$ and $p\geq 2$, $Q_{0}(n)>0$, implying $\mathbb{E}[\kappa(X_{n})]=\infty$ for every fixed $n\geq 1$ and $p\geq 2$. This completes the proof.
@@
Proposition 11 immediately implies that for every $r>1$ and $n\geq 1$ and $p\geq 2$,
$\mathbb{E}\!\left[\left(\frac{\sqrt{n}}{s_{\min}(X_{n})}\right)^{r}\right]=\infty,\qquad\mathbb{E}\!\left[\kappa(X_{n})^{r}\right]=\infty,$